OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

OpenStudy (michele_laino):

please, you have to calculate the first derivative of s(t), so what is: \[\frac{ ds }{ dt }=...\]

OpenStudy (anonymous):

is the answer -6?

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok. just changed the -6t to x and did d/dx(-6x) and got -6 and the -2 is 0

OpenStudy (anonymous):

can i ask another question? @Michele_Laino Use graphs and tables to find the limit and identify any vertical asymptotes of limit of 1 divided by the quantity x minus 7 squared as x approaches 7.

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

or 1/(x-7)^2

OpenStudy (anonymous):

finding the limit and and vertical asymptotes. i got 1 as the limit and 7 as any vertical asymptote

OpenStudy (anonymous):

is that correct?

OpenStudy (michele_laino):

vertical asymptote x=7 is right! limit when x is going to?

OpenStudy (anonymous):

x->7

OpenStudy (anonymous):

\[x\]

OpenStudy (anonymous):

\[x \rightarrow 7\]

OpenStudy (michele_laino):

we can only calculate this limit: \[\lim _{x \rightarrow 7+}\frac{ 1 }{ \sqrt{x-7} }=+ \infty\] since domain of your function is the set of all real x such that: x>7

OpenStudy (michele_laino):

namely limit when x goes to 7 from right!

OpenStudy (anonymous):

ok. thanks for the help. i'll give you a medal and fan you if you want.

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

that sounds wrong. *i will become a fan

OpenStudy (michele_laino):

ok!

hartnn (hartnn):

whats wrong in fanning someone ? O.o :P